OpenStudy (anonymous):

What is the solution to the rational equation x over x squared minus 9 plus 1 over x plus 3 equals 1 over 4x minus 12 ?

Nnesha (nnesha):

\[\huge\rm \frac{ x }{ x^2-9}+\frac{ 1 }{ x+3}=\frac{ 1 }{ 4x-12 }\] like this ?

OpenStudy (anonymous):

yes

Nnesha (nnesha):

alright first solve left side apply difference of square method to factor x^2-9 \[\rm a^2-b^2=(a+b)(a-b)\]

Nnesha (nnesha):

take square of both terms write ur answer as (sqrt of 1st term + sqrt of 2nd term) (sqrt of 1st term - sqrt of 2nd term) here is an example \[\rm x^2-16=(x+4)(x-4)\]

OpenStudy (anonymous):

x = -9 over 7 x = 15 x = 15 over 7 x = 9

OpenStudy (anonymous):

these are the answers i have

Nnesha (nnesha):

ok. so what are the factors of x^2-9 ?

OpenStudy (anonymous):

(x-9)(x+9)

Nnesha (nnesha):

hmm not what is the square root of 9 ?

OpenStudy (anonymous):

3

Nnesha (nnesha):